A manufacturer has 8 motors and 5 switches. In how many ways can you select 2 motors and 2 switches to build a project?
Answer by TimothyLamb(4379)

m: 8*7 = 56
s: 5*4 = 20
m+s: 56 + 20 = 76
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php